### v1

# Let Real Images be as a Judger, Spotting Fake Images Synthesized with Generative Models

Abstract
In the last few years, generative models have shown their powerful capabilities in synthesizing realistic images in both quality and diversity (i.e., facial images, and natural subjects). Unfortunately, the artifact patterns in fake images synthesized by different generative models are inconsistent, leading to the failure of previous research that relied on spotting subtle differences between real and fake. In our preliminary experiments, we find that the artifacts in fake images always change with the development of the generative model, while natural images exhibit stable statistical properties. In this paper, we employ natural traces shared only by real images as an additional predictive target in the detector. Specifically, the natural traces are learned from the wild real images and we introduce extended supervised contrastive learning to bring them closer to real images and further away from fake ones. This motivates the detector to make decisions based on the proximity of images to the natural traces. To conduct a comprehensive experiment, we built a high-quality and diverse dataset that includes generative models comprising 6 GAN and 6 diffusion models, to evaluate the effectiveness in generalizing unknown forgery techniques and robustness in surviving different transformations. Experimental results show that our proposed method gives 96.1% mAP significantly outperforms the baselines. Extensive experiments conducted on the widely recognized platform Midjourney reveal that our proposed method achieves an accuracy exceeding 78.4%, underscoring its practicality for real-world application deployment. The source code and partial self-built dataset are available in supplementary material.
1 Introduction
With the rapid development and maturity of generative models, the increasing proliferation of fake images has attracted widespread attention. Compared to Generative Adversarial Networks (GANs), diffusion models (DMs), as today’s SOTA generative models, exhibit better generation quality Dhariwal and Nichol (2021) and even support powerful text-to-image models such as DALL-E2 Ramesh et al. (2022), Stable Diffusion Rombach et al. (2022). Currently, one can use different types of generative models to create realistic faces or complex scene images and it is foreseeable that more generative models for image synthesis will emerge in the future. Therefore, it is the goal for the community to develop a more practical method to distinguish fake images that are synthesized with unknown forgery techniques as the unseen generative models will emerge inadvertently.
Figure 1: As the specific artifacts disappear, can fake detectors still work? For example, the detector trained on dog images generated by GANs is unable to detect ones generated by diffusion models.
The prior paradigm for fake image detection is to learn artifacts of fake images by capturing the subtle differences between real and fake images  Zhang et al. (2019); Wang et al. (2020). As shown in Figure 1, the detector distinguishes images as real or fake based on the presence of specific artifacts in the image. It is challenging to generalize to other unknown generative models, which may carry unseen or even imperceptible artifacts. Thus, a practical fake image detector should have strong capabilities in tackling fake images manipulated by unknown generative models. Unfortunately, the existing studies are trapped in the endless efforts to spot the artifacts by investigating the subtle differences between real and fake Corvi et al. (2023). The artifacts commonly explored are ad hoc to generative models and not applicable to diverse forgery techniques. As a defender, fake image detection seems to be perpetually behind the generation, only able to regulate obsolete fake images and helpless against the rapidly updating ones. Such a mouse-and-cat game reminds us whether the paradigm is trapped by exploring the subtle differences between real and fake, which is an endless effort as new forgery methods appear rapidly and the exploited artifacts for forensics will be invalid soon.
We argue that the challenges in the traditional paradigm arise from classifiers’ tendency to easily capture perceptible artifacts in training data. As these artifacts evolve or even vanish with generative model iterations, classifiers fail to detect new fake images. Merely investigating the flaws in fake images proves insufficient, given the constant emergence of unknown fake images that defenders have not seen. Instead, we consider a novel insight that the focus for fake image detection should be on the inherent similarities within real images, as opposed to the differences between real and fake. Our novel insight is motivated by the premise that real-world images possess an intrinsic naturalness that remains stable, whereas fake images continually evolve with technological innovation. We conjecture that there should be some shared features in real images, akin to that in fake images, which we term the natural trace.
Despite the stable statistical properties of natural images, precisely defining and capturing the shared feature of natural traces remains a significant challenge. In this work, we introduce Natural Trace Forensics (NTF) and adopt an alternative strategy, substituting the shared feature with homogeneous features, which are universally present in natural images. We develop a self-supervised feature mapping process to decouple real image representations, capturing stable homogeneous features accurately. Then, we integrate these features as an extra predictive target in parallel with binary classification, prompting the detector to aggregate these features closer to real images and separate them from fake ones. Guided by this strategy, the detector reduces its reliance on specific artifact patterns and gains the capability to detect images synthesized by unknown generative models.
To better evaluate whether our proposed method can be well generalized to other unknown generative models, we build a generated image dataset consisting of 12 SOTA generative models, including 6 GANs and 6 DMs. Our dataset covers a variety of categories, such as faces and scenes, to evaluate detectors’ capability across various types of fake images. In addition, we also evaluate the performance of our method in identifying images generated by Midjourney, which is currently a popular commercial tool for text-to-image generation. Experimental results show that our proposed method could discriminate GAN-based, DM-based, and Multi-step fake images (synthesized with at least two different generated models) in high confidence with an average accuracy of more than 96.1% and is sufficiently robust to various image perturbation transformations.
Our main contributions are summarized as follows:
  • We are the first to propose training fake image detection with natural traces, rather than merely a mix of real and fake images; our results also confirm its ability to detect different generative models and its potential as a universal judge for distinguishing fake images.
  • To conduct a comprehensive evaluation, we build a dataset including GAN-based, DM-based, and Multi-step manipulation for generating fake images. For the first time, we generate Multi-step fake images by employing multiple synthesis methods. This self-built dataset will be released to the community to advance the community to develop more effective and practical methods for distinguishing unseen forgery techniques.
  • Experimental results show the effectiveness of our proposed method in tackling fake images generated by SOTA GANs and diffusion models, giving an average precision of more than 96.18%, significantly outperforming the baselines. Additionally, our method is robust against the common image transformations and identifies the undistinguishable Midjourney in high confidence.
2 Related Work
2.1 Fake Images Synthesis
In recent years, generative models represented by GANs and diffusion models have been evolving towards maturity. GANs train a generator and a discriminator to compete with each other until the generator can produce realistic images, while DMs utilize diffusion and denoising Markov chains for training, surpassing GANs in training stability, diversity, and quality of generation Mazé and Ahmed (2023); Stypułkowski et al. (2024). This marks an evolution from producing or manipulating facial appearances, commonly termed DeepFakes Juefei-Xu et al. (2022), to generating contextually relevant text or image prompts. What’s more, with the emergence of new generative models, the artifacts of fake images are varied and more imperceptible, which also leads to the inability to detect fake images with low-level artifacts. It is foreseeable that the vast majority of artifact-based detection will fail when the fake image artifacts vanish. This progression has escalated public apprehension regarding the challenges in identifying fake images. In this paper, we are dedicated to developing a practical method for detecting fake images that are manipulated by unknown generative models.
2.2 Fake Images Detection
Existing research focuses on exploring subtle differences between real and fake images, and these can be categorized into explicit and implicit artifact-based methods.
Explicit-based. Some researchers noticed that they often contained specific artifacts or unnatural patternsZhang et al. (2019); Liu et al. (2023). Additionally, several studies focused on exploring GAN-based artifacts in the frequency domainFrank et al. (2020) and the failure to accurately re-enact certain biological features when generating fake facesHu et al. (2021); Tan et al. (2023b). These findings have motivated researchers to use explicit artifacts to detect fake images through simple classifiers. However, as generative models are continuously updated and iterated, these artifacts become imperceptible or even disappear, making detection methods reliant on explicit artifacts less effective.
Implicit-based. Wang et al.Wang et al. (2020) demonstrated that with appropriate training data and data augmentation, neural networks could detect other GAN-based images, while Tan et al. (2023a) used CNNs to transform images into gradient form to present a broader range of artifacts. These studies indicate that images generated by both GAN and diffusion models possess distinct “fingerprints" different from the real imagesYu et al. (2019); Sha et al. (2023). Nevertheless, there is still insufficient evidence to prove that generative models from different families have the same fingerprints that can be used for universal detection.
In this work, we propose an innovative method: training with real images for fake detection. It’s noteworthy that  Liu et al. (2022a) has similarities with ours, where they detect fake by analyzing noise patterns in real images. However, their method relies on the noise patterns of various devices. Our method uses natural traces extracted from real images for training. Compared to noise patterns, which may vary with different devices, can be complex to analyze, and are susceptible to intentional modification for evasion, natural traces consistently present in real images offer greater robustness.
Family Type Method Year Image Source # Images
Unconditional ProGAN 2017 CelebA-HQ 4.0k
StyleGAN2 2019 CelebA-HQ/FFHQ/LSUN 12.0k
GAN- ProjectedGAN 2021 FFHQ/LSUN/Landscape 12.0k
based VQGAN 2020 CelebA-HQ 4.0k
Diff-StyleGAN2 2022 FFHQ/LSUN 8.0k
Image-to-Image SimSwap 2020 CelebA-HQ 4.0k
Unconditional DDPM 2020 CelebA-HQ/LSUN 8.0k
DDIM 2021 CelebA-HQ 4.0k
DM- PNDM 2022 CelebA-HQ/LSUN 12.0k
based Image-to-Image DiffFace 2022 CelebA-HQ 2.0k
Prompt-guided LDM 2022 CelebA-HQ/LAION 8.0k
StableDM 2022 LAION 4.0k
GAN-GAN SimSwap_Style2 2024 CelebA-HQ 2.0k
SimSwap_VQ 2024 CelebA-HQ 2.0k
Multi- GAN-DM SimSwap_LDM 2024 CelebA-HQ 2.0k
step DiffFace_Style2 2024 CelebA-HQ 2.0k
DiffFace_Proj 2024 FFHQ 2.0k
DM-DM DiffFace_LDM 2024 CelebA-HQ 2.0k
Table 1: Statistics of the self-built dataset, including GAN-based, DM-based, and Multi-step synthesis.
3 A Diverse Generated Image Dataset
Due to the lack of DM-based generated images in current datasets, we create a dataset with a wide range of generative models. This dataset aims to enhance the evaluation of fake detection methods’ capability. It includes fake images generated by various models, alongside an equal number of real images from corresponding training sets for each method.
Our dataset covers two major families: GAN and Diffusion Models, with each fake image synthesized using a single generative model. Particularly, we have developed a novel multi-step fake image generation method, involving collaboration between two or more generative models, to achieve identity swapping or attribute editing between real and fake faces. To ensure diversity, the dataset comprises various categories of generation methods: unconditional generation, image-to-image generation, and prompt-guided generation, as shown in Table 1. Moreover, our carefully selected generative models exhibit fundamental differences in their generations, ensuring extensive representation of the dataset.
GAN-based forgery: We select six representative GANs to generate forgery images including unconditional and image-to-image. The innovation of ProGAN Karras et al. (2017) in introducing progressive training has a significant impact on subsequent research, while StyleGAN2 Karras et al. (2020) refines style control through decoupling image features and generates more realistic images. In terms of architectural improvements, ProjectedGAN Sauer et al. (2021) enhances generator feedback using pre-trained weights, while VQGAN Esser et al. (2021) and Diff-StyleGAN2 Wang et al. (2022) innovatively replace the backbone network with transformer and diffusion processes, respectively, resulting in higher image quality and more stable training processes. SimSwap’sChen et al. (2020) ID injection module achieves breakthroughs in arbitrary face swapping.
DM-based forgery: DM-based forgery surpasses GAN-based in terms of image quality and diversity. Here, we select six different DMs for creating fake images including unconditional, image-to-image, and prompt-guided. DDPMHo et al. (2020), as the initial diffusion model, lays the groundwork, with DDIMSong et al. (2020) and PNDMLiu et al. (2022b) enhancing execution speed and quality. DiffFaceKim et al. (2022) is the first identity-conditioned DDPM that uses diffusion models for face swapping. And LDMRombach et al. (2022), which employs pre-trained self-encoders to map pixels to latent space, combined with context learning in cross-attention layers for prompt-guided image generation, along with Stable Diffusion (StableDM), a popular LDM-based prompt-guided model, represent further advancements.
Multi-step: In the real scenario, the creator tends to employ multiple forgery techniques to achieve better forgery. We create a Multi-step face synthesis dataset where two face-swapping methods (SimSwap and DiffFace) based on GAN or diffusion models swap real faces onto synthetic ones, including three hybrid modes: GAN-GAN, GAN-DM, and DM-DM. This generation method provides simulated threats of artifact disappearance or blending in real-world scenarios.
In summary, our dataset includes high-resolution images ranging from $256\times 256$ to $1024\times 1024$ resolution. What’s more, diverse image categories support a more comprehensive evaluation of fake detection generalization. The dataset will be released to the community to advance this interesting research field.
Figure 2: NTF architecture. We first decouple the feature representation of real images into homogeneous and heterogeneous features. Next, the homogeneous features $z^{hom}$ participate in extended supervised contrastive training with the real and fake image features. Detector classifies real and fake images guided by the target of intra-class aggregation and inter-class separation. Better view in color.
Method ProGAN StyleGAN2 VQGAN ProjectedGAN Diff-StyleGAN2 SimSwap Average
AP $\uparrow$ ACC $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR
XceptionNet 99.31 84.47 0 31.05 100 99.95 0.10 0 100 99.92 0.15 0 100 99.10 1.80 0 97.89 84.03 0.20 31.75 99.85 90.13 19.74 0 99.51 92.93 3.66 10.47
Wang2020 97.88 85.33 27.05 2.00 86.15 56.33 87.30 0.05 61.21 50.52 98.90 0.05 62.18 50.35 99.20 0.10 87.79 68.45 61.30 1.80 98.36 65.04 69.78 0.05 82.26 62.67 73.92 0.68
Grag2021 100 99.98 0 0.05 100 100 0 0 99.88 91.62 16.75 0.05 99.72 94.62 10.75 0 99.82 94.47 11.05 0 94.12 50.49 98.90 0 98.92 88.53 22.91 0.02
Ojha2023 98.83 99.05 1.45 0.45 83.05 87.00 15.95 10.05 83.56 87.30 16.05 9.35 75.20 78.72 33.20 9.35 64.57 64.95 69.30 0.80 80.30 83.83 23.45 8.90 80.92 83.48 26.57 6.48
Self-Blended* 68.65 66.02 39.81 28.16 65.45 60.19 41.75 37.86 56.88 46.15 76.92 30.77 59.87 50.00 53.85 46.15 58.49 54.85 58.25 32.04 55.33 53.88 33.10 59.14 60.78 55.18 41.78 41.95
CADDM* 51.51 53.36 40.63 52.58 52.29 54.41 46.84 44.34 52.62 54.52 50.65 40.28 54.29 57.63 43.00 41.74 51.26 52.90 44.15 50.03 46.64 36.58 82.90 43.63 52.09 53.31 48.36 44.98
NTF 100 92.53 0 14.85 100 92.53 0 14.95 100 92.55 0 14.90 99.77 92.20 0.65 14.95 99.94 92.65 0.15 14.55 99.77 92.16 0.60 15.10 99.91 92.44 0.23 14.88
Table 2: Intra-family generalization on GANs. Performance of NFT and baselines in spotting 6 GAN-based generated images. These baselines include the classic approach, XcetionNet, Wang2020, and Grag2021, as well as the latest methods, Ojha2023, and CADDM. The method with *** means that the baseline is only evaluated on non-face images. The Average column represents the weighted average of the corresponding metrics. Among those, the best and second-best performances are highlighted in bold and underlined, respectively.
Method DDPM DDIM PNDM LDM StableDM DiffFace Average
AP $\uparrow$ ACC $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR
XceptionNet 74.69 72.37 5.07 50.19 75.78 73.89 12.36 39.86 89.61 49.75 99.95 0.61 65.75 65.02 8.80 61.16 41.07 40.25 28.47 91.03 59.23 40.77 99.45 19.01 66.21 57.01 14.04 59.03
Wang2020 58.15 50.15 99.65 0.05 60.90 50.15 99.60 0.10 66.35 49.95 99.55 0.05 67.87 50.40 99.20 0 39.81 49.48 99.55 1.50 34.75 49.75 99.79 2.40 54.64 49.98 99.56 0.68
Grag2021 79.92 50.12 99.75 0 83.70 50.25 99.45 0.05 86.46 50.10 99.31 0 99.98 98.78 2.45 0 98.17 63.55 72.90 0 63.19 54.35 91.29 0 85.24 61.19 77.53 0.01
Ojha2023 64.18 67.03 57.40 8.55 71.78 75.30 39.65 9.75 61.66 64.42 62.20 8.95 80.20 83.88 22.85 9.40 55.60 55.95 87.30 0.80 88.41 88.72 21.95 0.62 70.31 72.55 48.58 6.34
Self-Blended* 63.12 64.00 37.50 34.50 71.10 69.45 15.53 47.57 64.15 59.22 48.54 33.01 83.75 75.73 14.56 33.98 56.97 55.83 59.22 29.13 70.59 56.91 27.98 28.42 68.28 63.52 33.89 34.44
CADDM* 53.04 55.15 46.47 43.21 55.76 59.77 35.84 44.62 51.75 52.45 51.92 43.10 52.30 53.92 48.28 43.85 51.58 54.51 38.25 52.46 53.02 52.62 43.92 44.15 52.91 54.74 44.11 45.23
NTF 76.93 68.77 47.15 15.30 91.79 84.05 16.40 15.50 79.40 71.02 42.72 15.10 99.98 92.65 0.05 14.65 99.54 92.10 1.05 14.75 99.31 87.43 0.62 15.45 91.16 82.67 18.00 15.13
Table 3: Cross-family generalization on diffusion models. Performance of NFT and baselines in spotting 6 DM-based generated images.
Method SimSwap_Style2 SimSwap_VQ SimSwap_LDM DiffFace_Style2 DiffFace_Proj DiffFace_LDM Average
AP $\uparrow$ ACC $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR AP ACC FPR FNR
XceptionNet 44.60 50.00 100 0 46.56 50.00 100 0 48.51 50.00 100 0 100 100 0 0 100 100 0 0 37.97 50.00 100 0 62.94 66.67 66.67 0
Wang2020 89.32 55.7 88.60 0 76.07 50.80 98.20 0.20 79.38 50.00 98.4 0 69.89 78.91 90.88 0.05 65.19 50.10 99.40 0.40 54.68 50.00 100 0 72.42 55.92 95.91 0.11
Grag2021 99.76 85.70 28.60 0 97.07 61.90 76.00 0 95.24 55.40 89.20 0 100 99.85 0.66 0 80.78 79.40 41.20 0 53.03 50.00 100 0 87.65 72.04 55.94 0
Ojha2023 88.92 91.60 6.60 10.20 90.29 94.10 2.60 9.20 89.27 93.40 3.00 10.20 88.21 92.12 6.14 9.62 87.21 92.00 3.60 12.40 80.96 83.59 26.56 6.25 87.48 91.14 8.08 9.65
Self-Blended* 62.74 69.23 38.46 23.08 74.53 65.49 30.97 38.05 47.38 48.23 74.34 28.20 74.59 69.91 24.78 35.40 65.24 58.33 43.97 45.15 63.04 58.15 54.11 28.80 64.59 61.56 44.44 33.11
CADDM* 46.84 37.30 83.60 41.46 44.56 50.28 60.25 41.14 48.12 44.66 65.80 44.72 18.38 51.11 81.16 40.71 23.48 50.55 67.22 42.71 53.30 55.56 53.74 41.60 39.11 48.24 68.63 42.06
NTF 99.80 91.80 0.40 16.00 99.99 92.20 3.32 15.40 99.87 92.00 0.20 15.80 99.98 87.82 0 15.68 99.66 91.80 0.80 15.60 84.61 74.30 35.60 15.80 97.32 88.32 6.72 15.74
Table 4: Cross-family generalization on multi-step methods. Performance of NTF and baselines in spotting 6 multi-step generated images.
4 Our Method
In this work, to achieve a universal fake image detection and avoid the classifier’s reliance on specific artifact patterns, we propose a novel method, named Natural Trace Forensics (NTF), which involves training the classifier using the natural traces shared merely by real images as an additional predictive target. Figure 2 overviews the pipeline of NTF. We start by learning natural trace representations from real datasets. Then, under an extended supervised contrastive learning framework, the network is trained to align natural traces closer to real images and further from fake ones. With such constraints, the network is motivated to detect fakes based on the distance between images and natural traces. Next, we elaborate on how to learn the natural traces and apply the extracted traces for detection.
4.1 Natural Trace Representation Learning
We first explore the natural traces in real images to provide learnable features for the next fake image identification. However, it is impossible to analyze every real image in existence to identify shared features. To address this, we employ an innovative strategy using the same intrinsic features found in real images as a substitute for shared features. These intrinsic features, known as homogeneous features, are derived from the inherent properties and statistical regularities of images, which are commonly present in real images. We develop a self-supervised feature mapping mechanism to extract the homogeneous features. This mechanism decouples the natural image features into homogeneous features and heterogeneous features, where the latter are associated with specific images. As opposed to direct embedding features, feature decoupling not only enhances generalization by ensuring homogeneous features more accurately capture commonalities across images, but also improves feature quality by eliminating noise and redundant information. Additionally, the network should access a large variety of natural images. This exposure enables the network to learn the patterns of feature coupling in various types of real images so that it can decouple features for unseen images.
Formulation. We assume access to a large-scale dataset of real images $D_{r}$. Sample $x\in D_{r}$ is a augmented sample from a image $x_{r}\in\mathbb{R}^{H\times W\times 3}$ (of height $H$, and width $W$). As shown in Figure 2(a), our architecture consists of a feature encoder followed by two projection heads that map real image embeddings into homogeneous and heterogeneous features.
Specifically, feature encoder, $E$, producing feature embedding $e=E(x_{r})$ from the inputs, then decouples the $e$ into homogeneous and heterogeneous features, $z^{hom}=f^{hom}(e)\in\mathbb{R}^{C}$ and $z^{het}=f^{het}(e)\in\mathbb{R}^{C}$ through two projection heads, $f^{hom}$ and $f^{het}$, respectively, where $C$ is the dimensionality of the features. Let $i\in I\equiv\{1...2N\}$ be the index of an arbitrary augmented sample, where there are a total of N samples in a batch, each with two random augmentations ($2N$). The self-supervised contrastive loss of heterogeneous feature representations in real images can be formulated as follows:
$\mathcal{L}_{het}=-\sum_{i\in I}\log\frac{exp(z^{het}_{i}\cdot z^{het}_{p}/\tau)}{\sum_{k\in K(i)}exp(z^{het}_{i}\cdot z^{het}_{k}/\tau)},$ (1)
where the index $i$ is called the anchor, the index $p$ is called the positive, $\tau$ is a temperature hyperparameter and $K(i)\equiv I\backslash\{i\}$, which contains all the augmented samples except $i$ in a batch. Furthermore, the loss of homogeneous feature representations in real images can be formulated as:
$\mathcal{L}_{hom}=\max\arg_{i\in I,k\in K^{\prime}(i)}\parallel z^{hom}_{k}-z^{hom}_{i}\parallel^{2}_{F},$ (2)
where $\parallel\cdot\parallel_{F}$ denotes the Frobenius norm, $K^{\prime}$ contains $22N-$ 2 augmented samples, i.e., target $i$ is compared only with samples from different sources.
Considering the potential issues of high feature coupling in the embedding space, we further exploit soft orthogonality to reduce information redundancy and dependencies between these homogeneous and heterogeneous features:
$\mathcal{L}_{ort}=\sum_{i\in I}\cos(z^{hom}_{i},z^{het}_{i}),$ (3)
Finally, we combine these constraints to form the natural trace representation learning loss:
$\mathcal{L}_{tra}=\mathcal{L}_{hom}+\mathcal{L}_{het}+\lambda\mathcal{L}_{ort},$ (4)
where $\lambda$ is a scaling factor.
4.2 Fake Image Detection
To capture the homogeneous features as an additional predictive target, we employ an extended supervised contrastive learning to simultaneously encode real and fake images further for fake image detection. Specifically, we incorporate these homogeneous features into extended supervised contrastive learning and constrain their distance from positive and negative samples. This will motivate the detector to identify fake images based on distance. Note that the feature encoder is frozen during this stage.
Formulation. We now assume access a full dataset, $D=D_{r}\bigcup D_{f}$, where $D_{r}$ is used in the previous stage, $D_{f}$ is a dataset of fake images. Our architecture consists of a feature encoder, an auxiliary projection head for extended supervised contrastive learning, and another classification head for supervised classification(shown in Figure 2(b). With the auxiliary projection head, the real and fake feature embeddings are mapped to $z^{r}_{i}$ and $z^{f}_{i}$, respectively. To motivate the network to focus on the intra-class aggregation of real images more than the inter-class differences between real and fake images, we adopt the homogeneous features from the previous stage as extra positive instances for real anchors and additional negative instances for fake anchors. Formally, the extended supervised contrastive loss is as follows:
$\mathcal{L}_{ext}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\frac{exp(z_{i}\cdot z_{p}/\tau)}{\sum_{k\in K(i)}exp(z_{i}\cdot z_{k}/\tau)},$ (5)
where $P(i)$ is the set of all positive samples for the anchor $i$, $|P(i)|$ is its cardinality. Note that for real anchor, the auxiliary homogeneous feature $z^{hom}\in P(i)$, and vice versa, $z^{hom}\in K(i)$. To achieve fake image detection, for a given sample and label, the classifier $D$ is optimized on the binary cross entropy loss:
$\mathcal{L}_{ce}=-\frac{1}{N}\sum_{i}^{N}y_{i}\cdot\log(\hat{y_{i}})+(1-y_{i})\cdot\log(1-\hat{y_{i}}),$ (6)
Finally, the discriminative loss is given by:
$\mathcal{L}_{d}=\mathcal{L}_{ext}+\gamma\mathcal{L}_{ce},$ (7)
where $\gamma$ is a balance factor. Please refer to the technical appendix for more implementation details.
Figure 3: Ablation study on NTF architecture. All detectors were trained using the ProGAN and tested on other generative models. The designs of NTF architecture improve generalization ability. The red dotted line depicts chance performance.
Figure 4: Ablation study on training data. All detectors trained on different data sources (ProGAN or LDM) or different numbers of classes of the ProGAN data source (20 classes, 12 classes, and 4 classes). The diversity of training data improves generalization ability.
Method Wang2020 Grag2021 Ojha2023 NTF
Midjourney/Acc(%) $\uparrow$ 63.02 57.88 56.82 78.41
Table 5: Evaluation on commercial generative models. NTF shows 78.41% Acc on Midjourney, outperforming other baselines.
5 Experiments
5.1 Experiments Setup
Dataset. We use the dataset provided by Wang et al. (2020), which consists of 720K images for training and 4K images for validation. The fake images were generated by ProGANKarras et al. (2017), while an equal number of real images were sourced from the LSUNYu et al. (2015). We also conducted evaluation experiments on the self-built dataset, which covered GAN-based, DM-based, and multi-step fake images.
Evaluation Metrics. In evaluating the performance in spotting fake images synthesized with diverse generative models, we adopt four popular metrics to get a comprehensive result of our proposed method. Specifically, we report ACC (accuracy), AP (average precision), FPR (false positive rate), and FNR (false negative rate), respectively.
Baselines. For better evaluation, we compare with six baselines, including fake detection methods based on explicit artifacts and implicit artifacts. In experiments, we compare with the following six baselines: 1) XceptionNet Rossler et al. (2019) is widely employed as the baseline in the studies of DeepFake forensics; 2) Wang2020Wang et al. (2020) focuses on the artifacts exposed by CNN-generated images; 3) Grag2021Gragnaniello et al. (2021) uses spectral super-resolution to reconstruct visual cues for detection. 4) Ojha2023Ojha et al. (2023) uses a feature space not explicitly trained to distinguish real from fake images. 5) Self-BlendedShiohara and Yamasaki (2022) mixes image pairs with various masks to generate training data. 6) CADDMDong et al. (2023) focuses on local information so that the network ignores identity information leakage caused by irregular face changes. Except that Self-Blended and CADDM are trained on FaceForensics++ Rossler et al. (2019), the training set for the other comparison methods is consistent with ours. Note that Self-Blended and CADDM are limited to DeepFake datasets, and for the sake of fairness, they are not included in the test of non-face data.
Implementation Details. We use ResNet50 pre-trained on ImageNet as the feature encoder. All the projection heads contain two layers of MLPs with an output dimension of 128. For all datasets, we use a $224\times 224$ crop for both training and testing (random crop for training and center crop for testing). For extended supervised contrastive learning, we perform a random crop of the input image to 32 px. The optimizer is an SGD with a momentum of 0.9, an initial learning rate of 0.1, and an attenuation of 0.001. In the first stage, training is conducted for 200 epochs, followed by 10 epochs in the second stage. The empirical setting for $\lambda$, $\gamma$ is set at 0.1, 0.5.
Figure 5: Robustness to four image processing operations, i.e., Gaussian blur (a), JPEG compression (b), Gaussian Noise (c), Scaling (d).
Figure 6: The Grad-CAM++ visualization of the self-built dataset.
5.2 Effectiveness Evaluation
We explore the capability of NTF in detecting fake images from different generative models under the aforementioned experimental setup. Table 2-4 highlights its advanced performance through comparison with baselines.
Performance on GAN-based fake images. Experimental results in Table 2 demonstrate that NTF shows its capability in tackling unknown GAN-based fake images with the highest AP 99.91% and the lowest FPR 0.23% on average. In particular, NTF attains optimal detection performance on GAN-based fake images, which could be attributed to the similarity in the image generation principle of these models to that of ProGAN. That is precisely why certain baselines, such as XceptionNetRossler et al. (2019) and Grag2021Gragnaniello et al. (2021), have achieved high-performance generalization on GAN-based generative models. This indicates that existing fake image detection methods are adept at handling generalization scenarios within the same model family.
Performance on DM-based fake images. As illustrated in Table 3, NTF exhibits superior cross-family generalization capability on the DMs with the highest AP 91.16% and ACC 82.67% on average. Specifically, NTF improves the AP and ACC by nearly 5.92% and 10.12% compared to the best baseline. The detection accuracy of NTF surpasses all baselines across four DMs. However, it remains comparable to Grag2021Gragnaniello et al. (2021) on LDM and to XeceptionNetRossler et al. (2019) on DDPM. Overall, NTF detects different DM-base generation models in a more balanced way, although it is slightly inferior to some baselines on certain models.
Performance on multi-step fake images. As shown in Table 4, NTF detects fake images synthesized by six multi-step methods with the highest mAP of 97.32%. It is noteworthy that the method from Ojha2023Ojha et al. (2023) demonstrated an average ACC of 91.14%, slightly outperforming NTF. This can be attributed to their use of the ViT-L/14Dosovitskiy et al. (2020), a vision transformer variant pre-trained on CLIPRadford et al. (2021) for the backbone, aiding in effectively modeling details for real and fake image classification.
Performance on Challenging Midjourney. To better evaluate the performance in tackling commercial generative models, we assess the detection capability of NTF on fake images generated by Midjourney in Table 5. Experimental results show that our proposed method NTF gives an accuracy more than 78% which significantly outperforms the baselines.
In summary, NTF effectively detects GAN-based unknown generative models with a mAP 99.91%. It also demonstrates the capability to detect DM-based and multi-step generation methods with mAPs of 91.16% and 97.32% respectively, achieving an overall mAP of 96.18% across all datasets. Additionally, NTF shows a slightly higher FNR, which can be attributed to the inadequate coverage of real-world images in the training dataset.
5.3 Ablation Studies
We evaluated the effect of architecture and training data on the generalization ability of NTF. It initially employed homogeneous, heterogeneous, and soft orthogonality losses with Eq (4) for learning natural trace representations and it trains with the ProGAN/LSUN dataset consisting of 20 classes.
Effect of network architecture. We conducted experiments with different variants of NTF, exploring the following configurations: 1) without orthogonality loss, and 2) without heterogeneity and orthogonality loss. For each, We maintain ProGAN real/fake image data as training data. Figure 3 shows the performance of these variants on the same models. We find that the architectural design of NTF has an important role in generalization, improving performance on conditional generative models such as LDM, StableDM, and multi-step synthesis methods. This improvement is likely attributed to the implementation of soft orthogonality and heterogeneous constraints on upper bounds, which enables the NTF to learn homogeneous features capable of effectively separating traces of conditional generative models from natural traces. Moreover, this design also improves performance on unconditional generative models, particularly more significant in models like Diff-StyleGAN and PNDM.
Effect of training data Next, we investigate how the source and diversity of the training data influence a detector’s generalization ability. Our approach involves altering either the data source or the number of classes in the training set. Specifically, we train multiple detectors by 1) employing a pre-trained LDM, substituting the ProGAN, and 2) utilizing a subset of the full ProGAN dataset, excluding real and fake images from certain LSUN classes, as shown in Figure 4. With access only to the LDM dataset, the model displays impressive generalization capabilities, even though the dataset consists of 200k reals from LAIONSchuhmann et al. (2022) and 200k fakes generated by LDM. As expected, diversifying the training set does enhance generalization to some extent, but the benefits diminish with increasing diversity. This suggests the potential existence of a real image dataset capable of extracting universally present natural traces.
5.4 Robustness Evaluation
In this section, we mainly explore the robustness of our proposed method in surviving diverse input transformations, such as blur, compression, noise, and scaling.
As shown in Figure 5, NTF is typically robust to blur, compression, and noise operations. The performance of Wang2020Wang et al. (2020) is more consistent under different levels of blur/compression/noise. This is reasonable, as it employs random compression and blur for data augmentation during training. Although Ojha2023Ojha et al. (2023) also employs data augmentation, it also performs in addition to being less robust in blur operations. In addition, both NTF and other baselines show significant performance degradation in image scaling operations. When the image scaling factor is 0.2, the image content becomes imperceptible, making robustness in such extreme scenarios less critical. Please refer to the technical appendix for more details.
5.5 Qualitative Analysis
To understand how the network generalizes to different synthesis methods, we visualize the model saliency map. We apply GradCAM++ Chattopadhay et al. (2018) to NTF on the self-built dataset to visualize where models are paying their attention to images, as shown in Figure 6. This demonstrates how the network captures artifacts from different generative models; for example, with the StyleGAN2-generated image of a dog, the network focuses on the inconsistency in shadow.
6 Conclusion
In this paper, we escape the trap of exploring subtle differences between real and fake for fake image detection. Motivated by the presence of common shared features in real images, We propose a novel framework named NTF, pre-trained by natural trace representation learning and extended supervised contrastive learning, to significantly improve the generalization ability of fake image detection. Extensive experiments on the self-built dataset demonstrate that our method exhibits state-of-the-art generalization capability for unknown generative models. Our research also offers a fresh perspective on fake image detection, focusing on exploring stable detectable features rather than those that continuously change, thereby paving the way for future studies in this field.
7 Technical Appendix
In the technical appendix, we present the details of the self-built dataset, evaluation of the test set from Wang et al. (2020), some visualizations, and additional experiments.